# Noisy Quantum Learning Theory

### v1

@@ Abstract
We develop a framework for learning from noisy quantum experiments, focusing on fault-tolerant devices accessing uncharacterized systems through noisy couplings. Our starting point is the complexity class NBQP (“noisy BQP”), modeling noisy fault-tolerant quantum computers that cannot, in general, error-correct the oracle systems they query. Using this class, we show that for natural oracle problems, noise can eliminate exponential quantum learning advantages of ideal noiseless learners while preserving a superpolynomial gap between NISQ and fault-tolerant devices. Beyond oracle separations, we study concrete noisy learning tasks. For purity testing, the exponential two-copy advantage collapses under a single application of local depolarizing noise. Nevertheless, we identify a setting motivated by AdS/CFT in which noise-resilient structure restores a quantum learning advantage in a noisy regime. We then analyze noisy Pauli shadow tomography, deriving lower bounds that characterize how instance size, quantum memory, and noise control sample complexity, and design algorithms with parametrically similar scalings. Together, our results show that the Bell-basis and SWAP-test primitives underlying most exponential quantum learning advantages are fundamentally fragile to noise unless the experimental system has latent noise-robust structure. Thus, realizing meaningful quantum advantages in future experiments will require understanding how noise-robust physical properties interface with available algorithmic techniques.
@@ 1 Introduction
Quantum learning theory has advanced rapidly in recent years, and one of its key successes is recasting experimental protocols as learning problems, which admits a rich toolkit for determining when and how quantum computation-enhanced experiments can outperform conventional ones [Aar07, HHJ+17, WPS+17, Aar18, AdW18, ACH+18, HKP20, CW20, CCHL22, HKP21, ACQ22, CCHL23, HTFS23]. From this perspective, partially uncharacterized experimental systems serve as oracles providing quantum data, and the experimentalist implements a learning protocol that reveals the system’s properties through controlled interactions. When a quantum computer is available, it can be coherently coupled to the system of interest, enabling quantum computation on transduced data. In idealized settings, such quantum computation-enhanced experiments can provide a provable exponential advantage over conventional experimental protocols [HKP21, CCHL22, HBC+22, ACQ22].
@@
However, for most known examples of quantum experimental advantage, the quantum computer and its coupling to the experimental system are assumed to be noiseless. While a noisy quantum computer can be error-corrected, the experimental system given to us by Nature does not come embedded into an error-correcting code. We are therefore in a setting where an error-corrected quantum computer can only access the experimental system via a noisy coupling – one that must also mediate between the logical encoding of the quantum computer and the bare physical degrees of freedom of the experimental system. Moreover, since the experimental system corresponds to a partially unknown state or channel, it cannot generally be embedded into an error-correcting code by the experimentalist [ABOIN96]. Noise thus fundamentally changes the character of quantum computation-enhanced experiments and may obviate many of their known advantages.
@@
In this paper, we develop a theory of noisy quantum learning: for several prominent examples of quantum computation-enhanced experiments, we establish when noise eliminates quantum advantages and when such advantages persist. Central to our framework is the complexity class NBQP, modeling fault-tolerant quantum computers that access experimental systems through noisy couplings. We show that for many canonical quantum learning problems, noise reduces NBQP to the power of conventional experiments, yet for problems with latent error-correcting structure superpolynomial advantages survive.
@@
Concretely, noisy quantum learning theory reveals two distinct mechanisms by which noise erodes ideal quantum advantages. First, nearly all known idealized separations rely on learning tasks governed by intrinsically complex, high-weight structures, including states or dynamics that resist succinct classical representations [CCHL22, HBC+22, NTGK25, CGY24, KGKB25]. For example, the quantum-enhanced sample-complexity gap in estimating physically motivated Pauli observables grows with the weight of the observables [CGY24], and classical simulation of quantum dynamics becomes intractable when high-weight Paulis dominate the evolution [Sch11, BC23]. Our results show, however, that precisely these high-complexity components are the most vulnerable to noise. As a result, the very structures that underwrite ideal quantum advantages also provide the primary mechanism through which noise destroys them.
@@
Second, exponential quantum advantages achieved through multi-copy measurements typically require entanglement across a number of qubits that scales extensively with the problem size [HBC+22, CCHL22, GHYZ24]. Operationally, such algorithms are built from a small set of standard primitives, most notably maximally entangled (Bell-basis) measurements and many-qubit SWAP operations, that underlie idealized superpolynomial speedups in quantum learning, property testing, and computation [HM13, LMR14, BOW19, MdW18]. In our setting, implementing these primitives on noisy, uncharacterized quantum systems leads to an exponential blow-up in sample complexity. Absent additional structure in the physical system that can be exploited for error correction, the primitives that enable ideal quantum speedups cease to provide robust advantages in the noisy regime.
@@
Making these concepts rigorous requires ruling out adaptive learning protocols that might otherwise compensate for noise. Our framework for proving the inefficiency of noisy quantum learning protocols, even with adaptive strategies, builds on the learning tree formalism of [CCHL22], and in particular techniques from [CCHL23]. In the latter work, noisy quantum computers incapable of fault-tolerant error correction were formalized as the complexity class NISQ, and it was shown that there exist oracles $O$ for which $\textsf{NISQ}^{O}\subsetneq\textsf{BQP}^{O}$, meaning that even adaptive NISQ protocols are superpolynomially slower than noiseless quantum computers for certain learning problems. Our refinement is to introduce a more physically realistic complexity class, NBQP, corresponding to noisy quantum computers with a constant error rate per qubit per circuit layer. The noise rate is taken to be below the threshold for fault-tolerant error correction, so that $\textsf{NBQP}=\textsf{BQP}$. Accordingly, we refer to learning protocols captured by $\textsf{BQP}^{O}$ as ideal quantum learning, and those captured by $\textsf{NBQP}^{O}$ as natural quantum learning, which are depicted in Fig. 1. We demonstrate that for certain natural quantum learning problems, known to have an exponential advantage over conventional experiments with unentangled access, the corresponding oracles $O$ enforce $\textsf{NBQP}^{O}\subsetneq\textsf{BQP}^{O}$. A consequence is that noise can obviate the advantage of even adaptive quantum computing-enhanced experiments over conventional experiments.
@@
Given that several canonical results in quantum learning theory break down in noisy settings, what kinds of quantum advantages can still persist in the presence of such noise? A key insight is that many physical systems possess endogenous robustness: thermalizing systems have built-in noise resilience [CRL10, BCSB19, KKCG25], and long-range correlations in many-body quantum systems can be protected by renormalization group structure [PYHP15, KK17, LBC25]. The robustness of the physical system and the fault-tolerant architecture of the quantum computer can meet in the middle, allowing exponential quantum advantages to survive. We demonstrate this through illustrative examples in which a natural error-correcting structure present in a class of quantum states synergizes with the error correction of the quantum computer. Furthermore, we construct natural quantum learning problems for which $\textsf{NISQ}^{O}\subsetneq\textsf{NBQP}^{O}$, showing that error-correcting the quantum computer provides a superpolynomial advantage even when the coupling to the experimental system remains noisy.
@@ 2 Main Results
In Section 2.1, we formalize the class NBQP and prove two superpolynomial oracle separations: ${\textnormal{{NBQP}}}^{O_{1}}\subsetneq{\textnormal{{BQP}}}^{O_{1}}$ and ${\textnormal{{NISQ}}}^{O_{2}}\subsetneq{\textnormal{{NBQP}}}^{O_{2}}$. Next, we move from complexity separations to concrete noisy learning tasks. In Section 2.2, we study the NBQP complexity of testing the purity of a quantum state, showing that an ideal exponential quantum advantage is fundamentally obstructed by noise unless the class of states in question has a latent error-correcting structure. We give a concrete example of such a structure in a physically-motivated toy setting: assessing the purity of a black hole microstate in the bulk of a tensor-network model of holographic duality using only noisy measurements of the boundary state. In Section 2.3, we analyze Pauli shadow tomography, demonstrating a tradeoff between sample complexity, the noise rate, and the complexity of the estimated observables. While it is still possible to have a polynomial advantage of certain quantum-enhanced experiments vis-à-vis conventional experiments in the noisy setting, we find that noise exponentially degrades the most prominent idealized quantum advantages. Technical details of the proofs are deferred to the appendix.
@@ 2.1 Oracle separations between NBQP, BQP, and NISQ
NBQP contains all problems solvable in polynomial time by a noisy quantum computer with polynomial-size circuits, where all operations are subject to constant depolarizing noise per qubit at a rate below the threshold of known quantum error-correcting codes.
@@
Under the above definition, an NBQP quantum computer can implement error correction and perform recovery after each layer of gates on any register embedded into a code. The threshold theorem [ABO97] then implies $\textsf{NBQP}=\textsf{BQP}$. However, the situation changes when we introduce oracle access. To make this concrete, consider coupling a fault-tolerant quantum computer to a partially uncharacterized quantum material in the laboratory. Even if the quantum computer itself is fully error-corrected, the interaction with the material occurs at the physical-qubit level and is therefore noisy. We cannot simply transduce the material’s state into an error-correcting code; even if this were technologically feasible, the encoding procedure could corrupt the very quantum information we are trying to learn about. The resulting asymmetry, that errors on the computation can be corrected but errors on oracle queries cannot, opens the possibility of nontrivial separations between $\textsf{NBQP}^{O}$ and $\textsf{BQP}^{O}$. While error mitigation may be possible for particularly structured oracles, we demonstrate that the inability to perform true error correction on the unknown system can starkly degrade our ability to learn from experiments, leading to superpolynomial oracle separations between NBQP and BQP.
@@
To prove this separation, we adapt the lifted Simon oracle construction of [CCHL23]. Conceptually, the standard Simon oracle is modified so that slight perturbations to its input result in completely uninformative outputs, making the oracle highly nonrobust to NBQP-type noise. More concretely, given a function $f:\{0,1\}^{n}\to\{0,1\}^{n}$, consider the usual Simon’s promise: either (i) $f$ is injective, or (ii) there exists a nonzero $s\in\{0,1\}^{n}$ such that $f(x)=f(x\oplus s)$ for all $x$ (a Simon’s function). The lift $\tilde{f}$ acts on $2n$ input bits but is nontrivial only when the last $n$ bits are $0^{n}$. A noiseless quantum computer can simply restrict its queries to strings of the form $x\,0^{n}$ and then run the standard Simon’s algorithm [Sim94] on the first $n$ bits. Thus the promise problem “is $f$ injective or a Simon’s function?” lies in ${\textnormal{{BQP}}}^{O_{1}}$ with $O(n)$ queries.
For ${\textnormal{{NBQP}}}^{O_{1}}$, oracle calls act on physical qubits, and in our model each such call is preceded and followed by a layer of depolarizing noise. Even if the algorithm tries to prepare “good” queries of the form $x\,0^{n}$, these noise layers quickly flip some of the trailing $n$ bits, so with high probability the actual query lies outside the special subspace on which $\tilde{f}$ encodes $f$; on those inputs $O_{1}$ simply outputs 0. In effect, to a noisy learner $O_{1}$ is almost indistinguishable from a trivial oracle, and exponentially many queries in $n$ are required for such a learner to tell the difference. Using the hybrid/distinguishability framework of [CCHL23], namely their channel-level hybrid lemma and a node-perturbation argument in our learning-tree model (formalized as Lemmas D.16 and D.17), we show that any $\lambda$-noisy circuit making $N$ oracle calls has output distribution within $N^{2}e^{-\Omega(\lambda n)}$ total variation distance of the distribution obtained by replacing every oracle call with the identity channel. Under this identity oracle the two promise cases (injective versus Simon’s $f$) induce exactly the same distribution, so any ${\textnormal{{NBQP}}}^{O_{1}}$ algorithm with $N=\mathrm{poly}(n)$ cannot achieve constant distinguishing advantage, whereas a ${\textnormal{{BQP}}}^{O_{1}}$ algorithm can. This yields the strict separation ${\textnormal{{NBQP}}}^{O_{1}}\subsetneq{\textnormal{{BQP}}}^{O_{1}}$ claimed in the theorem.
Despite this limitation, an NBQP machine can implement polynomial-depth fault-tolerant quantum circuits, which are believed to be strictly more powerful than the logarithmic-depth circuits achievable by noisy intermediate-scale (NISQ) devices [FGHZ05, HS05, HFK+22, ACC+23]. Leveraging this gap in coherent depth, we prove a superpolynomial oracle separation between NISQ and NBQP, showing that fault-tolerant learners retain an advantage over near-term devices even when all access to the experimental system is noisy.
@@
For the proof of this theorem, we explicitly leverage the fundamental gap between NISQ and NBQP machines: the latter has ability to perform quantum error correction, and thus can execute deep quantum circuits. Following [CCL23], we start from the $d$-level Shuffling Simon’s Problem, a variant of Simon’s problem challenging for shallow quantum circuits to solve. Here, a Simon function $f$ on $n$ bits is embedded and randomly permuted inside a much larger domain so that only a hidden subset of inputs carries any information about the secret $s$. We then define an encoded shuffling oracle $\mathcal{O}^{\mathrm{enc}}_{f,d}$ that acts like this shuffled Simon oracle on a logical code space of a fixed fault-tolerant quantum error correction scheme, and trivially outside it. By Theorem 4.11 of [CCL23], there is a noiseless depth-$O(d)$ circuit that recovers $s$, and an NBQP machine can simulate this circuit fault-tolerantly (for the noise rate $\lambda$ below threshold) with only polynomial overhead, and so $\textsf{Enc-}d\textsf{-SSP}\in{\textnormal{{NBQP}}}^{O_{2}}$.
@@
The lower bound against NISQ proceeds in two steps. First, building on the analysis of [CCL23], we show that any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm has exponentially small success probability on the encoded problem: even if it makes polynomially many depth-$d$ quantum queries, the domain in which the Simon’s function is embedded is so large that the algorithm is very unlikely to query the relevant subspace. While classical advice between bounded-depth subroutines can reveal successively more information about this “hidden domain”, the number of permutations applied to the Simon’s function is, by construction, too large for any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm to locate the domain. Mathematically, we carry out this argument by demonstrating that, with high probability, all oracle queries made by the algorithm can be replaced by “shadow” queries that agree with the true oracle outside a small wrapper set containing the hidden domain, but output no information on the domain itself; even after this swap, the output distribution of the algorithm is hardly altered. Upon this replacement, the algorithm learns nothing about $s$, as it no longer has access to the embedded Simon’s function (Lemma D.11), and can thus do no better than random guessing. This establishes an exponentially small success probability for any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm.
@@
Second, we show that any NISQ algorithm making polynomially many oracle calls can be simulated, up to vanishing total variation distance, by such a $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm: we cut each noisy circuit at a fixed depth threshold and use KL-divergence bounds from [CCHL23] to argue that replacing deeper noisy circuits by shallow ones only changes the leaf distribution of the learning tree by $o(1)$. Combining these two ingredients and applying Le Cam’s two-point method, we obtain that no NISQ algorithm with $\mathrm{poly}(n)$ queries can recover the Simon’s secret with success probability at least $2/3$, whereas an NBQP algorithm can, yielding the oracle separation $\textsf{NISQ}^{O_{2}}\subsetneq\textsf{NBQP}^{O_{2}}$.

### v2

@@ Abstract
We develop a framework for learning from noisy quantum experiments in which fault-tolerant devices access uncharacterized systems through noisy couplings. Introducing the complexity class NBQP (“noisy BQP”), we model noisy fault-tolerant quantum computers that cannot generally error-correct the oracle systems they query. Using this class, we prove that while noise can eliminate the exponential quantum learning advantages of unphysical, noiseless learners, a superpolynomial gap remains between NISQ and fault-tolerant devices. Turning to canonical learning tasks in noisy settings, we find that the exponential two-copy advantage for purity testing collapses under local depolarizing noise. Nevertheless, we identify a setting motivated by AdS/CFT in which noise-resilient physical structure restores this quantum learning advantage. We then analyze noisy Pauli shadow tomography, deriving lower bounds characterizing how instance size, quantum memory and noise jointly control sample complexity, and design algorithms with parametrically matching scalings. We study similar tradeoffs in quantum metrology, and show that the Heisenberg-limited sensitivity of existing error-correction-based protocols persists only up to a timescale inverse-polynomial in the error rate per probe qubit. Together, our results demonstrate that the primitives underlying quantum-enhanced experiments are fundamentally fragile to noise, and that realizing meaningful quantum advantages in future experiments will require interfacing noise-robust physical properties with available algorithmic techniques.
@@ 1 Introduction
Quantum learning theory has advanced rapidly in recent years, and one of its key successes is recasting experimental protocols as learning problems, which admits a rich toolkit for determining when and how quantum computation-enhanced experiments can outperform conventional ones [Aaronson2006Learnability, HaahHarrowJiWuYu2017SampleOptimal, Wang_2017, Aaronson2018ShadowTomography, ArunachalamDeWolf2018QuantumSampleComplexity, Aaronson_2019, classical_shadow, cotler2020quantum, chen2021exponentialseparationslearningquantum, Huang_2021, Aharonov_Cotler_Qi_2022, nisq, Huang_2023]. From this perspective, partially uncharacterized experimental systems serve as oracles providing quantum data, and the experimentalist implements a learning protocol that reveals the system’s properties through controlled interactions. When a quantum computer is available, it can be coherently coupled to the system of interest, enabling quantum computation on transduced data. In idealized settings, such quantum computation-enhanced experiments can provide a provable exponential advantage over conventional experimental protocols [Huang_2021, chen2021exponentialseparationslearningquantum, Huang_2022, Aharonov_Cotler_Qi_2022].
@@
However, for most known examples of quantum experimental advantage, the quantum computer and its coupling to the experimental system are assumed to be noiseless. While a noisy quantum computer can be error-corrected, the experimental system given to us by Nature does not come embedded into an error-correcting code. We are therefore in a setting where an error-corrected quantum computer can only access the experimental system via a noisy coupling – one that must also mediate between the logical encoding of the quantum computer and the bare physical degrees of freedom of the experimental system. Moreover, since the experimental system corresponds to a partially unknown state or channel, it cannot generally be embedded into an error-correcting code by the experimentalist [aharonov1996limitationsnoisyreversiblecomputation]. Noise thus fundamentally changes the character of quantum computation-enhanced experiments and may obviate many of their known advantages.
@@
In this paper, we develop a theory of noisy quantum learning: for several prominent examples of quantum computation-enhanced experiments, we establish when noise eliminates quantum advantages and when such advantages persist.
@@
Concretely, noisy quantum learning theory reveals two distinct mechanisms by which noise erodes ideal quantum advantages. First, nearly all known idealized separations rely on learning tasks governed by intrinsically complex, high-weight structures, including states or dynamics that resist succinct classical representations [chen2021exponentialseparationslearningquantum, Huang_2022, nöller2025infinitehierarchymulticopyquantum, weiyuan_paulis, King_2025]. For example, the quantum-enhanced sample-complexity gap in estimating physically motivated Pauli observables grows with the weight of the observables [weiyuan_paulis], and classical simulation of quantum dynamics becomes intractable when high-weight Paulis dominate the evolution [Schollw_ck_2011, begušić2023fastclassicalsimulationevidence]. Our results show, however, that precisely these high-complexity components are the most vulnerable to noise. As a result, the very structures that underwrite ideal quantum advantages also provide the primary mechanism through which noise destroys them.
@@
Second, exponential quantum advantages achieved through multi-copy measurements typically require entanglement across a number of qubits that scales extensively with the problem size [Huang_2022, chen2021exponentialseparationslearningquantum, gong2024samplecomplexitypurityinner]. Operationally, such algorithms are built from a small set of standard primitives, most notably maximally entangled (Bell-basis) measurements and many-qubit SWAP operations, that underlie idealized superpolynomial speedups in quantum learning, property testing, and computation [Harrow:2012gwf, Lloyd_2014, bădescu2017quantumstatecertification, montanaro2018surveyquantumpropertytesting]. In our setting, implementing these primitives on noisy, uncharacterized quantum systems leads to an exponential blow-up in sample complexity. Absent additional structure in the physical system that can be exploited for error correction, the primitives that enable ideal quantum speedups cease to provide robust advantages in the noisy regime.
@@
Making these concepts rigorous requires ruling out adaptive learning protocols that might otherwise compensate for noise. Our framework for proving the inefficiency of noisy quantum learning protocols, even with adaptive strategies, builds on the learning tree formalism of [chen2021exponentialseparationslearningquantum], and in particular techniques from [nisq]. In the latter work, noisy quantum computers incapable of fault-tolerant error correction were formalized as the complexity class NISQ, and it was shown that there exist oracles $O$ for which $\textsf{NISQ}^{O}\subsetneq\textsf{BQP}^{O}$, meaning that even adaptive NISQ protocols are superpolynomially slower than noiseless quantum computers for certain learning problems. Our refinement is to introduce a more physically realistic complexity class, NBQP, corresponding to noisy quantum computers with a constant error rate per qubit per circuit layer. The noise rate is taken to be below the threshold for fault-tolerant error correction, so that $\textsf{NBQP}=\textsf{BQP}$. Accordingly, we refer to learning protocols captured by $\textsf{BQP}^{O}$ as ideal quantum learning, and those captured by $\textsf{NBQP}^{O}$ as natural quantum learning, which are depicted in Fig. 1. We demonstrate that for certain natural quantum learning problems, known to have an exponential advantage over conventional experiments with unentangled access, the corresponding oracles $O$ enforce $\textsf{NBQP}^{O}\subsetneq\textsf{BQP}^{O}$. A consequence is that noise can obviate the advantage of even adaptive quantum computing-enhanced experiments over conventional experiments.
@@
Given that several canonical results in quantum learning theory break down in noisy settings, what kinds of quantum advantages can still persist in the presence of such noise? Our work suggests two avenues to realize practically meaningful gains in noisy quantum-enhanced experiments. Many physical systems possess endogenous robustness: thermalizing systems have built-in noise resilience [Chesi_2010, Brandao_2019, Kastoryano_2025], and long-range correlations in many-body quantum systems can be protected by renormalization group structure [happy_2015, Kim_2017, Lake_2025]. A useful insight is that the robustness of the physical system and the fault-tolerant architecture of the quantum computer can meet in the middle, allowing exponential quantum advantages to survive. We demonstrate this through illustrative examples in which a natural error-correcting structure present in a class of quantum states synergizes with the error correction of the quantum computer. Furthermore, we construct natural quantum learning problems for which $\textsf{NISQ}^{O}\subsetneq\textsf{NBQP}^{O}$, showing that error-correcting the quantum computer provides a superpolynomial advantage even when the coupling to the experimental system remains noisy.
Even in noisy regimes where asymptotic exponential quantum advantages disappear [Chen_2024_pauli, seif2024, Oh_2024, Liu_2025], sub-exponential advantages can persist. Analyzing noisy quantum-enhacned experimental protocols requires tracking how complexity depends jointly on noise rates, quantum resources, and finite instance size, since practical experimental behavior is governed by these factors rather than asymptotic limits. Our Pauli-learning separation illustrates that even for a canonical quantum learning task, noise-sensitive effects yield meaningful quantum advantages that interpolate between idealized exponential separations and polynomial ones as the noise rate scales with instance size. Beyond quantum state learning, we apply noise- and resource-aware analyses to metrology, showing that single-parameter protocols achieving Heisenberg-limited sensitivity through active error correction can sustain this scaling only up to a timescale set by instance size, control speed, and device-probe coupling noise.
@@ 2 Main Results
In Section 2.1, we formalize the class NBQP and prove two superpolynomial oracle separations: ${\textnormal{{NBQP}}}^{O_{1}}\subsetneq{\textnormal{{BQP}}}^{O_{1}}$ and ${\textnormal{{NISQ}}}^{O_{2}}\subsetneq{\textnormal{{NBQP}}}^{O_{2}}$. Next, we move from complexity separations to concrete noisy learning tasks. In Section 2.2, we study the NBQP complexity of testing the purity of a quantum state, showing that an ideal exponential quantum advantage is fundamentally obstructed by noise unless the class of states in question has a latent error-correcting structure. We give a concrete example of such a structure in a physically-motivated toy setting: assessing the purity of a black hole microstate in the bulk of a tensor-network model of holographic duality using only noisy measurements of the boundary state. In Section 2.3, we analyze Pauli shadow tomography, demonstrating a tradeoff between sample complexity, the noise rate, and the complexity of the estimated observables. Using similar techniques, we show that error-correction-based metrology protocols retain Heisenberg-limited sensitivity only up to a sensing timescale that scales inverse-polynomially with sensor dimension and noise rate. While it is still possible to have a meaningful polynomial advantage of certain quantum-enhanced experiments vis-à-vis conventional experiments in the noisy setting, we find that noise exponentially degrades the most prominent idealized quantum advantages. Technical details of the proofs are deferred to the appendix.
@@ 2.1 Oracle separations between NBQP, BQP, and NISQ
NBQP contains all problems solvable in polynomial time by a noisy quantum computer with polynomial-size circuits, where all operations are subject to constant depolarizing noise per qubit at a rate below the threshold of known quantum fault-tolerance schemes.
@@
Under the above definition, an NBQP quantum computer can implement error correction and perform recovery after each layer of gates on any register embedded into a code. The threshold theorem [aharonov1999faulttolerantquantumcomputationconstant] then implies $\textsf{NBQP}=\textsf{BQP}$. However, the situation changes when we introduce oracle access. To make this concrete, consider coupling a fault-tolerant quantum computer to a partially uncharacterized quantum material in the laboratory. Even if the quantum computer itself is fully error-corrected, the interaction with the material occurs at the physical-qubit level and is therefore noisy. We cannot simply transduce the material’s state into an error-correcting code; even if this were technologically feasible, the encoding procedure could corrupt the very quantum information we are trying to learn about. The resulting asymmetry — that errors on the computation can be corrected but errors on oracle queries cannot — opens the possibility of nontrivial separations between $\textsf{NBQP}^{O}$ and $\textsf{BQP}^{O}$. While error mitigation may be possible for particularly structured oracles, we demonstrate that the inability to perform true error correction on the unknown system can starkly degrade our ability to learn from experiments, leading to superpolynomial oracle separations between NBQP and BQP.
@@
To prove this separation, we adapt the lifted Simon oracle construction of [nisq]. Conceptually, the standard Simon oracle is modified so that slight perturbations to its input result in completely uninformative outputs, making the oracle highly nonrobust to NBQP-type noise. More concretely, given a function $f:\{0,1\}^{n}\to\{0,1\}^{n}$, consider the usual Simon’s promise: either (i) $f$ is injective, or (ii) there exists a nonzero $s\in\{0,1\}^{n}$ such that $f(x)=f(x\oplus s)$ for all $x$ (a Simon’s function). The lift $\tilde{f}$ acts on $2n$ input bits but is nontrivial only when the last $n$ bits are $0^{n}$. A noiseless quantum computer can simply restrict its queries to strings of the form $x\,0^{n}$ and then run the standard Simon’s algorithm [Simon] on the first $n$ bits. Thus the promise problem “is $f$ injective or a Simon’s function?” lies in ${\textnormal{{BQP}}}^{O_{1}}$ with $O(n)$ queries.
For ${\textnormal{{NBQP}}}^{O_{1}}$, oracle calls act on physical qubits, and in our model each such call is preceded and followed by a layer of depolarizing noise. Even if the algorithm tries to prepare “good” queries of the form $x\,0^{n}$, these noise layers quickly flip some of the trailing $n$ bits, so with high probability the actual query lies outside the special subspace on which $\tilde{f}$ encodes $f$; on those inputs $O_{1}$ simply outputs 0. In effect, to a noisy learner $O_{1}$ is almost indistinguishable from a trivial oracle, and exponentially many queries in $n$ are required for such a learner to tell the difference. Using the hybrid/distinguishability framework of [nisq], namely their channel-level hybrid lemma and a node-perturbation argument in our learning-tree model (formalized as Lemmas D.16 and D.17), we show that any $\lambda$-noisy circuit making $N$ oracle calls has output distribution within $N^{2}e^{-\Omega(\lambda n)}$ total variation distance of the distribution obtained by replacing every oracle call with the identity channel. Under this identity oracle the two promise cases (injective versus Simon’s $f$) induce exactly the same distribution, so any ${\textnormal{{NBQP}}}^{O_{1}}$ algorithm with $N=\mathrm{poly}(n)$ cannot achieve constant distinguishing advantage, whereas a ${\textnormal{{BQP}}}^{O_{1}}$ algorithm can. This yields the strict separation ${\textnormal{{NBQP}}}^{O_{1}}\subsetneq{\textnormal{{BQP}}}^{O_{1}}$ claimed in the theorem.
Despite this limitation, an NBQP machine can implement polynomial-depth fault-tolerant quantum circuits, which are believed to be strictly more powerful than the logarithmic-depth circuits achievable by noisy intermediate-scale (NISQ) devices [fenner2003boundspowerconstantdepthquantum, fanout, Haferkamp_2022, arora2023quantum]. Leveraging this gap in coherent depth, we prove a superpolynomial oracle separation between NISQ and NBQP, showing that fault-tolerant learners retain an advantage over near-term devices even when all access to the experimental system is noisy.
@@
For the proof of this theorem, we explicitly leverage the fundamental gap between NISQ and NBQP machines: the latter has ability to perform quantum error correction, and thus can execute deep quantum circuits. Following [Chia_2023], we start from the $d$-level Shuffling Simon’s Problem, a variant of Simon’s problem challenging for shallow quantum circuits to solve. Here, a Simon function $f$ on $n$ bits is embedded and randomly permuted inside a much larger domain so that only a hidden subset of inputs carries any information about the secret $s$. We then define an encoded shuffling oracle $\mathcal{O}^{\mathrm{enc}}_{f,d}$ that acts like this shuffled Simon oracle on a logical code space of a fixed fault-tolerant quantum error correction scheme, and trivially outside it. By Theorem 4.11 of [Chia_2023], there is a noiseless depth-$O(d)$ circuit that recovers $s$, and an NBQP machine can simulate this circuit fault-tolerantly (for the noise rate $\lambda$ below threshold) with only polynomial overhead, and so $\textsf{Enc-}d\textsf{-SSP}\in{\textnormal{{NBQP}}}^{O_{2}}$.
@@
The lower bound against NISQ proceeds in two steps. First, building on the analysis of [Chia_2023], we show that any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm has exponentially small success probability on the encoded problem: even if it makes polynomially many depth-$d$ quantum queries, the domain in which the Simon’s function is embedded is so large that the algorithm is very unlikely to query the relevant subspace. While classical advice between bounded-depth subroutines can reveal successively more information about this “hidden domain”, the number of permutations applied to the Simon’s function is, by construction, too large for any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm to locate the domain. Mathematically, we carry out this argument by demonstrating that, with high probability, all oracle queries made by the algorithm can be replaced by “shadow” queries that agree with the true oracle outside a small wrapper set containing the hidden domain, but output no information on the domain itself; even after this swap, the output distribution of the algorithm is hardly altered. Upon this replacement, the algorithm learns nothing about $s$, as it no longer has access to the embedded Simon’s function (Lemma D.11), and can thus do no better than random guessing. This establishes an exponentially small success probability for any $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm.
@@
Second, we show that any NISQ algorithm making polynomially many oracle calls can be simulated, up to vanishing total variation distance, by such a $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm: we cut each noisy circuit at a fixed depth threshold and use KL-divergence bounds from [nisq] to argue that replacing deeper noisy circuits by shallow ones only changes the leaf distribution of the learning tree by $o(1)$. Combining these two ingredients and applying Le Cam’s two-point method, we obtain that no NISQ algorithm with $\mathrm{poly}(n)$ queries can recover the Simon’s secret with success probability at least $2/3$, whereas an NBQP algorithm can, yielding the oracle separation $\textsf{NISQ}^{O_{2}}\subsetneq\textsf{NBQP}^{O_{2}}$.